# RealMath: A Continuous Benchmark for Evaluating Language Models on Research-Level Mathematics

### v1

@@ A.1 Prompts for Judge Models
17- We DON’T want the theorems that state "X $\in$\$ complexity class Y" since Y can belong to a bigger complexity class Z, so the answer is not unique.
@@ A.3 Manual Review
For the datasets we analyze in Section 4, we rely on a human-based review to ensure the quality of the samples. As noted in Section 3.2, for Math.arXiv we removed 5.2% bad samples, whereas for CS.arXiv we discarded 7.5% samples. Figure 8 shows the percentages of manually discarded items that fall into each category (35 samples in total for Math.arXiv and 9 samples for CS.arXiv). These samples were either theorems or QA pairs that did not have a unique answer but which LLMs failed to filter out in the previous stages of the automated pipeline.
@@ A.5 Pipeline for Mathematics Stack Exchange
The pipeline we describe in Section 3 is designed for arXiv papers. To retrieve and process samples from Mathematics Stack Exchange, we need to adapt some steps. We first queried the corresponding API to retrieve users’ questions with tags such as [limits], [definite-integrals], [integration]—which were expected to have a clear numerical answer—as well as the top-ranked answer, if available. We extracted the HTML-formatted text and used an LLM to process the user’s post into a statement and proof format, mirroring the theorems extracted from arXiv papers. We then applied a manual filtering step to ensure statements have a fixed answer and that the answer is indeed correct (for this, we checked cited sources, as well as the top-rated answer posted by other users). Finally, we fed the formulated theorem to an LLM to generate a QA pair and filter trivial samples.
@@ B.1 Finetuning
In Figure 10, we use OpenAI’s fine-tuning service to fine-tune GPT-4o-mini on our dataset. We present the detailed training and test loss in Figure 11. To ensure reliability, we repeat the evaluation five times and report the mean and standard deviation of the results. Surprisingly, we observe no statistically significant improvement.

### v2

@@ A.1 Prompts for Judge Models
17- We DON’T want the theorems that state "X $\in$$ complexity class Y" since Y can belong to a bigger complexity class Z, so the answer is not unique.
@@ A.3 Manual Review
For the datasets we analyze in Section˜4, we rely on a human-based review to ensure the quality of the samples. As noted in Section˜3.2, for Math.arXiv we removed 5.2% bad samples, whereas for CS.arXiv we discarded 7.5% samples. Figure 8 shows the percentages of manually discarded items that fall into each category (35 samples in total for Math.arXiv and 9 samples for CS.arXiv). These samples were either theorems or QA pairs that did not have a unique answer but which LLMs failed to filter out in the previous stages of the automated pipeline.
@@ A.5 Pipeline for Mathematics Stack Exchange
The pipeline we describe in Section˜3 is designed for arXiv papers. To retrieve and process samples from Mathematics Stack Exchange, we need to adapt some steps. We first queried the corresponding API to retrieve users’ questions with tags such as [limits], [definite-integrals], [integration]—which were expected to have a clear numerical answer—as well as the top-ranked answer, if available. We extracted the HTML-formatted text and used an LLM to process the user’s post into a statement and proof format, mirroring the theorems extracted from arXiv papers. We then applied a manual filtering step to ensure statements have a fixed answer and that the answer is indeed correct (for this, we checked cited sources, as well as the top-rated answer posted by other users). Finally, we fed the formulated theorem to an LLM to generate a QA pair and filter trivial samples.
@@ B.1 Finetuning
In Figure˜10, we use OpenAI’s fine-tuning service to fine-tune GPT-4o-mini on our dataset. We present the detailed training and test loss in Figure˜11. To ensure reliability, we repeat the evaluation five times and report the mean and standard deviation of the results. Surprisingly, we observe no statistically significant improvement.